A Bibliography of Publications in American

Mathematical Monthly: 1960–1969

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.07

Title word cross-reference

#5375 [13]. #7672 [14]. 1915 [10]. 29 [13]. 32 [14]. 7 [13]. 72 [14]. 84c [13]. 90 [13]. ACM [13]. adjacent [3]. Amer [14, 13].

America [10]. American [10]. analysis [15]. Analytical [1]. Anniversary [10]. Association [10]. C [13]. Calculation [4]. Combinatorial [15]. Comm [13]. Commemoration [10]. Comment [13]. Computer [4]. computers [15]. consecutive [8]. December [10]. eigenvalues [13, 12]. elementary [1]. entries [13, 12]. Errata [14]. Error [4]. Estimation [4]. expressions [1]. factors [8]. Fiftieth [10]. function [7]. functions [1].

Generation [2, 13]. groups [6, 14].

inequality [9, 5]. integer [13, 12]. integers [8]. interchanges [3]. Issue [10].

J [13]. J.-C [13].

law [9, 5].

magic [11]. Math [14, 13]. Mathematical [10, 10]. Matrices [13, 12, 2]. Monthly [14, 13, 10]. MR [14, 13].

no [13].

parallelogram [9, 5]. Permutation [3]. primality [7]. prime [8]. Problem [9, 5]. Published [10].

Ramanujan [7]. Renaud [13].

Semitopological [6, 14]. similarity [2, 13]. Solutions [9]. squares [11].

tau [7]. tau-function [7]. test [2, 13]. transformations [2, 13]. triangle [9, 5].

Very [11].

References

Sholander:1960:AEE

[1] Marlow Sholander. Analytical expressions and elementary functions. American Mathematical Monthly, 67(??):213– 214, 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ortega:1964:GTM

[2] James M. Ortega. Generation of test matrices by similarity transformations. Communications of the ACM, 7(6):377– 378, June 1964. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic). See comment [13].

Lehmer:1965:PAI

[3] D. Lehmer. Permutation by adjacent interchanges. American Mathematical Monthly, 72(??):36–46, ???? 1965. CO-DEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ashenhurst:1965:EEC

[4] R. L. Ashenhurst and N. Metropolis. Error estimation in computer calculation. American Mathematical Monthly, 72(??): 47–58, ???? 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1965:PTI

[5] Donald E. Knuth. Problem 5264, the triangle inequality and the parallelogram law. American Mathematical Monthly, 72 (2):193, February 1965. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [9].

Anonymous:1965:SG

[6] Anonymous. Semitopological groups. American Mathematical Monthly, 72(??): 996–998, ???? 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See errata [14].

Lehmer:1965:PRT

[7] D. H. Lehmer. The primality of Ramanujan’s tau-function. American Mathematical Monthly, 72(2):15–18, ???? 1965. CO-DEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1965:PFC

[8] D. H. Lehmer. The prime factors of consecutive integers. American Mathematical Monthly, 72(2):19–20, ???? 1965. CO-DEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Buchman:1966:SPT

[9] E. O. Buchman and W. G. Dotson, Jr. Solutions to problem 5264, the triangle inequality and the parallelogram law. American Mathematical Monthly, 73(2): 211–212, February 1966. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [5].

Anonymous:1967:FAI

[10] Anonymous. Fiftieth anniversary issue of the American Mathematical Monthly. published in commemoration of the Mathematical Association of America on December 31, 1915. American Mathematical Monthly, 74(1 (Part 2)):1–110 (??), ???? 1967. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1968:VMS

[11] Donald E. Knuth. Very magic squares.

American Mathematical Monthly, 76:260– 264, March 1968. ISSN 0002-9890 (print), 1930-0972 (electronic).

Renaud:1983:MIE

[12] J.-C. Renaud. Matrices with integer entries and integer eigenvalues. American Mathematical Monthly, 90(3):202– 203, ???? 1983. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See comment [13].

Ortega:1985:CMI

[13] James M. Ortega. Comment on: “Matrices with integer entries and integer eigenvalues” by J.-C. Renaud [Amer. Math. Monthly 90 (1983), no. 3, 202–203, MR 84c:15024 ] and “Generation of test matrices by similarity transformations” by Ortega [Comm. ACM 7 (1964), 377–378, MR 29 #5375]. American Mathematical Monthly, 92(7):526, ???? 1985. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [2, 12].

Bohn:1987:ESG

[14] S. Elwood Bohn and Jong Lee. Errata: “Semitopological groups” [Amer. Math. Monthly 72 (1965), 996–998, MR 32 #7672]. American Mathematical Monthly, 94(10):994, ???? 1987. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [6].

Hall:1965:CAC

[15] Marshall Hall, Jr. and Donald E. Knuth. Combinatorial analysis and computers. American Mathematical Monthly Slaught Memorial Papers, 10:21–28, February 1965. Also numbered as vol. 72, no. 2, Part II.